1. A and B consume only two goods, cider (C) and dumplings (D). A has an initial endowment of 10 bottles of C and 30 of D. Bob has an initial endowment of 50 bottles of cider and 50 dumplings. Alice’s utility function is uA(CA,DA) = 9ln(CA) + 10ln(DA), where CA and DA represent consumption of C and D, respectively. B’s utility function is uB(CB,DB) = CBXDB, where CB and DB denote B's consumption of C and D.


a) Find the competitive equilibrium, i.e. the price ratio, of this exchange economy and the resulting equilibrium allocation.


b) Find the expression of the contract curve for this economy and use your answer to check that the equilibrium allocation you found in (b) is indeed Pareto optimal.



1
Expert's answer
2021-08-12T05:09:28-0400

Given A and B as consumers

C=cider

D=dumption

A=10c,30d

B=50c,50d

"a)\\c=50-10=40\\

d=50-30=20\\

40:20\\

2:1\\"

b)

"A
10c+50d\\B\\10c+30d"